Afternoon tasks: ---------------- DA and instrument checkout completed with no problems. Jon & Warren worked some more this afternoon on the secondary galil motor connections. They replaced pins and replaced spliced wires. Steph ran the secondary back & forth several times, but could not reproduce the "not stopping on a full step error". Let's hope this is fixed. We shall see. Robert & Steph verified that the same fiducial tables were used for both pointing models taken this dark run. The difference between the two models that was noted last night is not due to the fiducial tables. Pete & Craig called Russell to see if there were any TCC related problems that could have caused the difference. PT: --- Biases, dome and twilight flats taken. We opened at 20:30. Cart 1, no star with either model. Cart 3. field at ra=313, dec=-6. fk5infiber went to 260 +47? second fk5infiber went to nearby star. Offset ra+12, dec+2 to center star, then we got guide stars. Spectro 1 fill faults at 22:19 did not correspond to actual warming.